A study shows that the cost of raising a child from birth to age one is more than Php 92 000. A

random sample of 36 families reveal a mean of Php 95 000 with a standard deviation of Php 5 000.

Based on these sample data, can it be concluded that the study is correct in its claim? Use 0.01 level

of significance.


1
Expert's answer
2021-07-06T13:51:29-0400

"H_0: \mu ≤ 92000 \\

H_1: \mu > 92000 \\

n=36 \\

\bar{x}= 95000 \\

s= 5000 \\

α=0.01"

Critical value at 0.01 significance level with df =35 is 2.43

"t = \frac{95000-92000}{5000 / \sqrt{36}}= 3.60"

Since test statistics t > 2.43, we have sufficient evidence to reject H0.

We conclude at 0.01 level that we have enough evidence to support the claim.